# Rising bubbles draw surface patterns: a numerical study

### v1

@@ Abstract
Small bubbles rising in a chain can self-organize into regular patterns upon reaching a liquid’s free surface. This phenomenon is investigated through direct numerical simulations. By varying the bubble release period, distinct branching patterns characterized by different numbers of arms are observed. These macroscopic regular configurations arise from localized non-contact repulsion and pair collisions between bubbles as they arrive at the free-surface emergence site. A theoretical model is proposed to quantitatively relate the number of branches to the bubble release period. The model also predicts probabilities of observing specific arm counts in reality. This study provides insights into broader nonlinear pattern formation and self-organization phenomena.
@@ I Introduction
When a bubble ascends to the gas-liquid interface, it may rapidly burst, accompanied by generation of small droplets[17, 18, 19, 20]. In cases where there are contaminants in the liquid, the bubble can float and persist for a considerable time[21]. The complexity of bubble-free surface interactions becomes more pronounced in systems that are specially designed and controlled[22, 23].
This study centers on the collective behavior of periodically-released small bubbles rising in a chain toward the free surface. After reaching the interface, the bubbles exhibit lateral migration within the horizontal plane, forming spatial patterns observed from above. While most of these patterns appear disordered, under specific conditions, the bubbles spontaneously self-organize into stable, regular configurations with rotational symmetry around the vertical ascent trajectory. An experimental demonstration of these regular surface patterns is shown in Fig. 1. In this experiment, air bubbles with a uniform diameter of approximately $0.8\,\mathrm{mm}$ are periodically injected from a syringe into a vessel partially filled with silicone oil. A spiral pattern is clearly observed on the free surface when looking from above. The density, dynamic viscosity, and surface tension coefficient of the silicon oil are $935\,\mathrm{kg/m^{3}}$, $0.01\,\mathrm{Pa{\cdot}s}$, and $0.02\,\mathrm{N/m}$, respectively. The bubble release interval is $16\,\mathrm{ms}$.
@@
Despite the likelihood that such phenomenon has been observed since very early times, the literature on it is limited. The first article on this specific topic appears to be Yoshikawa et al. [24], where the authors conducted an experimental investigation of the surface bubble patterns as an analogy to the regular leaf arrangement in the context of phyllotaxis study (see, e.g., Ref. [25]). A phenomenological model was later developed by the same research group[26]. These two papers acknowledge the importance of bubble-bubble interactions and flow advection. Nevertheless, the mechanisms by which these factors contribute to pattern formation require further elucidation. In this paper, we revisit these regular surface patterns “drawn” by rising bubbles from a numerical perspective, aiming to gain a deeper understanding of the essential underlying mechanism of this self-organized phenomenon.
The rest of the paper is arranged as follows: Section II describes the numerical setup of the simulations, followed by section III, where the results are presented and discussed. Section IV provides more theoretical considerations. A summary is given in section V.
@@ II Numerical setup
respectively, where $\bm{u}$ is the velocity vector, $p$ is pressure, $\bm{g}$ is the gravitational acceleration, and $\bm{f}_{\sigma}$ represents the surface tension effect. $\rho$ and $\mu$ are the effective density and dynamic viscosity, respectively. They are calculated as weighted average based on the local gas volume fraction $\alpha$:
@@
Eqs. (1) and (2) are solved using the second-order Bell-Colella-Glaz method[28]. The advection of the volume fraction is handled using the volume-of-fluid (VOF) method, with a piecewise linear interface calculation (PLIC) method[29] employed for interface reconstruction. The surface tension term is defined as $\bm{f}_{\sigma}=\sigma\kappa\nabla\alpha$, where $\sigma$ denotes the surface tension coefficient. The interfacial curvature, $\kappa$, is evaluated using the method of particles[30]. Additionally, an improved multi-marker technique[27] is employed in the VOF method to prevent bubble coalescence, ensuring that bubbles can float on the free surface without rupture in the simulations.
The accuracy of the solver is validated by simulating a single bubble rising in quiescent liquid. Four different liquids are tested with their physical parameters the same as those of the cases S3, S5, S6, and S8 in Ref. [6]. The Morton numbers, defined as $Mo=g\mu_{l}^{4}/(\rho_{l}\sigma^{3})$, are 0.11, $9\times 10^{-4}$, $1\times 10^{-4}$, and $9\times 10^{-7}$ for these four cases, respectively. For each liquid, the volume-equivalent spherical diameter of the bubble, $d$, is varied across a relatively wide range. The computational domain is a liquid-filled box with a Cartesian mesh composed of cubic cells. The cell size is set to approximately $d/10$, maintaining consistent spatial resolution for different bubble sizes.
@@
The computational domain for simulations of the surface patterns is a square box partially filled with liquid (see Fig. 3a). A Cartesian coordinate system is established with the origin $O$ located at the geometric center of the box’s bottom wall. The $z$-axis is oriented upward, while the $x$- and $y$-axes align with the side walls of the box. The dimensions of the box are $0.1,0.1]\,\text{m}\times[-0.1,0.1]\,\text{m}\times[0,0.075]\,\text{m}$, with the undisturbed free surface located at z=0.055z=0.055\,m. Gas bubbles with a volume-equivalent diameter of $d=4$ mm are released periodically from a source at (0, 0, 0.005)(0,\,0,\,0.005) m. Free-slip conditions are imposed at the top and bottom boundaries of the computational domain, while periodic conditions are applied on all side boundaries. Since all the boundaries are reasonably far from the area of interest, i.e., the central region of the free surface, the influence of the boundary conditions on the bubble behavior is negligible. A uniform mesh with $N_{x}\times N_{y}\times N_{z}=512\times 512\times 192$ cubic cells is used, ensuring that there are about 10 mesh cells within the diameter of a spherical bubble.
Figure 3: (a) Illustration of the computational domain (bubble size not to scale). (b) Simulated bubbles rising in a chain, with the free surface (black) and streamlines (gray) shown in a cross-sectional view.
In all simulations, the gravitational acceleration is set to $g=10\,\mathrm{m/s^{2}}$ in the z-z direction. The liquid ($l$) and gas ($g$) phases have the following properties: $\rho_{l}=1000\,\mathrm{kg/m^{3}}$, $\rho_{g}=30\,\mathrm{kg/m^{3}}$, $\mu_{l}=2\times 10^{-2}\,\mathrm{Pa\cdot s}$, and $\mu_{g}=2\times 10^{-5}\,\mathrm{Pa\cdot s}$. The surface tension coefficient $\sigma=0.02\,\mathrm{N/m}$. These physical parameters are similar to those used by Yoshikawa et al. [24], allowing a reasonable comparison between the numerical results and the experimental data in Ref. [24]. Given the low density ratio ($\rho_{g}/\rho_{l}=0.03$) and viscosity ratio ($\mu_{g}/\mu_{l}=0.001$) between gas and liquid, the effects of gas phase properties are expected to be minimal. The Ohnesorge and Bond numbers are calculated as $Oh=\mu_{l}/\sqrt{\rho_{l}\sigma d}\approx 0.071$ and $Bo=\rho_{l}gd^{2}/\sigma=8$, respectively. The corresponding Morton number is $Mo=Oh^{4}Bo=2\times 10^{-4}$, lying between the values of cases S5 and S6 in Fig. 2. The bubbles exhibit moderate deformation during ascent. No path instabilities are observed with the parameters investigated.
@@
For the single-bubble cases shown in Fig. 2, a bubble acquires its terminal velocity soon after release. However, because of the wake effects from preceding bubbles, for a bubble rising in a bubble chain, a much larger height is required for its velocity to reach a constant value. In the simulations of the current study, the bubbles are still gradually accelerating before approaching the free surface. Consequently, the release depth $h$ (or its dimensionless form $h/d$) might also be a relevant parameter. Nevertheless, whether the bubbles reach the terminal velocity is unlikely to have any qualitative impact on the mechanism of the pattern formation.
@@ III.1 Overview
With the VOF method, the motion and deformation of each bubble are fully resolved. Fig. 3(b) shows a cross-sectional view of the numerical result for a typical case we have simulated. The bubbles rise along a straight path in a chain and are arrested by the free surface. As can be seen from the streamlines, the bubble chain induces a vertical liquid jet along the central axis. This jet later on converts into a divergent radial flow in a thin layer beneath the free surface. Owing to the overshooting effect of the bubbles and jet, a small circular bump forms on the free surface, centered at the bubble emergence site. These observations agree well with those reported in Ref. [26].
Four representative surface patterns observed in the simulations are presented in the top panels of Fig. 4. These patterns are obtained, from (a) to (d), by decreasing the bubble release period $T$ while keeping all other parameters fixed. Panel (a) shows a disordered mode, characterized by a random distribution of bubbles on the free surface. Panel (b) displays a two-armed mode, where the bubbles align along two branches in a straight line. Further decreasing the release period results in the multi-armed modes, as shown in panels (c) and (d). The corresponding movies for all these four cases can be found in the supplemental material of this paper.
Figure 4: (a) to (d): Four typical surface patterns observed in the numerical simulations. The dashed lines serve as visual guides to distinguish between different branches. The numbers indicate the sequence in which successive bubbles join these branches. (e) to (h): Divergence angle as a function of bubble release sequence number for the corresponding cases shown in the top row. The red dashed lines show the averaged values. From left to right, the release periods are 100, 50, 42, and 36 ms, respectively.
It is known that the pattern formation is closely related to the directions in which the bubbles are emitted in sequence. In the bottom panels of Fig. 4, the difference $\varphi$ (marked in Fig. 5b below, clockwise positive) in the emission direction between consecutive bubbles, referred to as the “divergence angle” in Ref. [24], is shown as a function of bubble release sequence number. As will be subsequently elucidated, $\varphi$ is fully determined by the local dynamics in the vicinity of the bubble emergence site. In detail, four major mechanisms are identified to be responsible for the formation of different surface patterns. They are (i) pair collisions, (ii) non-contact repulsion, (iii) bump confinement, and (iv) flow advection. The first two mechanisms both describe the interaction between bubbles. Pair collision occurs when two bubbles are close enough with their mass center distance $\ell\lesssim d$, and is associated with significant bubble squeezing, deformation, and quick velocity changes. Whereas non-contact repulsion is effective when $\ell\gtrsim d$ and is less violent.
@@ III.2 Disordered and fixed two-armed modes
The disordered mode is seemingly the most common situation. It occurs when the bubble release period $T$ is relatively large. This mode is characterized by random pair collisions between bubbles. For instance, as bubble 1-1 rises to the free surface, it collides with bubble 2-2 which is already at the emergence site. The two bubbles are knocked away by each other in opposite directions, creating a $180^{\circ}$ divergence angle. By the time bubble 0 reaches the surface, the influence of the previous collision has vanished. Bubble 0 stays at the center and collides with the subsequent bubble 1, resulting in opposite emission directions that do not necessarily align with those of bubbles 2-2 and 1-1. As a consequence, the divergence angle jumps between $180^{\circ}$ and other random values (Fig. 4e). Additionally, a large $T$ results in a relatively weak liquid jet, such that the bubbles tend to aggregate near the central region of the free surface. Both of these features hinder the formation of regular patterns with distinguishable numbers of arms.
The transition mechanism from the disordered mode to a fixed two-armed mode (Fig. 4b) is straightforward: When $T$ is decreased below a certain threshold value, bubble 1-1, after collision with bubble 2-2, does not have enough time to move far away from the center before bubble 0 rises to the free surface. Consequently, bubble 0 collides with bubble 1-1 and bounces towards the opposite direction to the latter’s position vector. Similarly, bubble 1 is emitted in the opposite direction to that of bubble 0, following the path of bubble 1-1. In this way, the bubbles alternatively join two branches with a constant divergence angle of $\varphi=180^{\circ}$, as depicted in Fig. 4(f).
@@ III.3 Multi-armed modes
The multi-armed modes occur when the bubble release period is further decreased. They exhibit richer phenomena compared with the disordered and fixed two-armed situations. Figs. 4(c) and 4(d) show a fixed five-armed pattern and a spiral three-armed pattern, respectively. A 3D view of a single bubble trajectory in the five-armed case is displayed in Fig. 5(a). The projection of this trajectory onto the horizontal plane is shown in Fig. 5(b) alongside the trajectories of two consecutive bubbles.
Figure 5: (a) Free surface (gray) and trajectory (cyan) of a single bubble in a five-armed case. (b) Trajectories of three consecutive bubbles projected onto the horizontal plane. The green dashed lines indicate the emission directions of bubbles 2-2 and 1-1.
@@
The twisted trajectories in Fig. 5 indicate complex bubble dynamics in the vicinity of the bubble emergence site. Yoshikawa et al. [24] note that, despite the arms may rotate, “the motion of each bubble is purely radial.” This is true except for in the short-period initial phase of the bubble motion, as can be seen in Fig. 5(b). In order to understand the entire process that determines the divergence angle, we divide the trajectory of a single bubble into six stages, labeled I to VI, in Fig. 6. The dynamics of the bubbles and their connection with the mechanisms listed in section III.1 are described in detail below.
@@
Firstly, a significant difference between multi-armed modes and the simpler fixed two-armed mode is that the former cannot be explained solely by binary bubble collisions. This adds complexity to the analysis of the underlying physics, necessitating the inclusion of mechanism (ii), i.e., non-contact repulsion. This mechanism is illustrated in Fig. 7, where three snapshots of the three-armed situation are shown in chronological order. During this stage, bubble 0 is still rising below the free surface without colliding with preceding bubbles. Floating on the free surface with a short distance from the center where the bubbles emerge, bubble 2-2 acts as an obstacle that partially blocks the rightward motion of the liquid flow. As a consequence of this asymmetry, bubble 0’s rising path is deflected to the left, as if it experiences repulsion from bubble 2-2. Since there is no direct contact between bubbles 0 and 2-2 throughout the interaction, the horizontal displacement of bubble 0 from the central axis is mild. We denote this displacement by $\bm{\lambda}_{1}$. This process corresponds to stage I in Fig. 6(a). Similar interactions also occur between bubbles 0 and 1-1. However, since bubble 1-1 is only slightly deviated from the center, its influence on bubble 0’s horizontal motion is much weaker than the influence of bubble 2-2. The repulsion from other floating bubbles is negligible due to their larger distance from bubble 0. Therefore, $\bm{\lambda}_{1}$ is roughly in the opposite direction of bubble 2-2’s position vector in the horizontal plane, as shown in the inset of Fig. 5(b).
@@
The motion of bubble 0 in stage II is governed by mechanism (i), i.e., pair collision. Fig. 8 depicts the collision process between bubbles 0 and 1-1 in the cross-section where the centers of the two bubbles are located. The snapshot in Fig. 8(a) is taken at about the same time as that in Fig. 7(c). When approaching the free surface, bubble 0 is impeded by bubble 1-1, squeezed and pushed away by it. During this process, bubble 0 acquires a horizontal displacement roughly in the opposite direction of bubble 1-1’s position vector, which can also be observed in the inset of Fig. 5(b). We denote this displacement by $\bm{\lambda}_{2}$.
@@
Although mechanisms (i) and (ii) both have effects to accelerate bubble 0, the bubble’s radial velocity falls back to nearly zero by the end of stage II. This can be seen from Fig. 9(a), where the radial velocities of bubbles 2-2, 1-1, 0, and 1 are shown as functions of time. The velocity variation can also be deduced from Fig. 6. In this figure, the dots in the trajectories are successively sampled from the numerical results with a constant time interval. Therefore, their sparseness is an indicator of the bubble’s horizontal velocity magnitude. This deceleration of bubble 0 is due to mechanism (iii) mentioned in section III.1, i.e., the confinement of the bump. As revealed in Fig. 8(c), after collision with bubble 1-1, bubble 0 is located in the very top region of the bump. Thus, its horizontal motion is restricted by the capillary effect of the deformed and more vertical side surface of the bump.
Figure 9: (a) Radial velocity versus time for bubbles 2-2, 1-1, 0, and 1 (from left to right). Stages are labeled with respect to bubble 0. (b) Time dependence of bubble 0’s radial (blue), vertical (green), and circumferential (pink, clockwise positive) velocities. The red line represents the distance between the bubble mass center and the undisturbed free surface. Time is set to zero at the instant of bubble release.
It is noteworthy that, up to this point, the bubble’s displacement from the central axis is no greater than half its radius. However, according to Fig. 6(a), the final emission direction (the red dash-dotted line) of bubble 0 has been essentially determined. All the remaining stages play a role to “launch” this nearly stationary bubble and transport it in the same direction as its current position vector.
The collision with bubble 0 also promotes bubble 1-1 to leave the central region. As a consequence, bubble 0 has more room to flatten, and the bump surface becomes less steep. This relieves the constraint imposed by the bump, allowing bubble 0 to regain a small velocity away from the center advected by the radial flow mentioned earlier. This process corresponds to stage III in Figs. 6 and 9(a).
@@
Being pushed away from the center by bubble 1, the subsequent motion of bubble 0 in stage V is driven by the divergent radial flow (mechanism (iv) in section III.1). This effect of flow advection is demonstrated by the streamlines shown in Fig. 10(a). The bubble continues moving within the bump until it encounters the edge of it, where the bubble experiences a temporary deceleration due to the transition between the bumped and horizontal free surface. This reason can be deduced from Fig. 9(b), where the vertical dashed line indicates that the local minimum radial velocity occurs exactly when the bubble reaches the flat liquid surface from a higher elevation.

### v2

@@ Abstract
Small bubbles rising in a chain can self-organize into regular patterns upon reaching a liquid’s free surface. This phenomenon is investigated through direct numerical simulations. By varying the bubble release period, distinct branching patterns characterized by different numbers of arms are observed. These macroscopic patterns are found to stem from localized interactions between the bubbles, the free surface, and the surrounding liquid flow. A heuristic model is proposed that quantitatively relates the number of arms to the bubble release period and predicts the probability of observing specific arm counts. This study provides insights into broader nonlinear pattern formation and self-organization phenomena.
@@ I Introduction
When a bubble ascends to the gas-liquid interface, it may rapidly burst, accompanied by generation of small droplets[17, 18, 19, 20]. In cases where liquid viscosity is large or when contaminants/surfactants are present, bubbles can float and persist for a considerable time[21, 22]. The complexity of bubble-free surface interactions becomes more pronounced in systems that are specially designed and controlled[23, 24].
This study centers on the collective behavior of periodically-released small bubbles rising in a chain toward the free surface. After reaching the interface, the bubbles exhibit lateral migration within the horizontal plane, forming spatial patterns observed from above. While most of these patterns appear disordered, under specific conditions, the bubbles spontaneously self-organize into stable, regular configurations with rotational symmetry around the vertical ascent trajectory. An experimental demonstration of these regular surface patterns is shown in Fig. 1. In this experiment, air bubbles with a uniform diameter of approximately $0.8\,\mathrm{mm}$ are periodically injected from a syringe into a vessel partially filled with silicone oil. A spiral pattern is clearly observed on the free surface when looking from above. The density, dynamic viscosity, and surface tension coefficient of the silicone oil are $935\,\mathrm{kg/m^{3}}$, $0.01\,\mathrm{Pa{\cdot}s}$, and $0.02\,\mathrm{N/m}$, respectively. The bubble release interval is $16\,\mathrm{ms}$.
@@
Despite the likelihood that such phenomenon has been observed since very early times, the literature on it is limited. The first article on this specific topic appears to be Yoshikawa et al. [25], where the authors conducted an experimental investigation of the surface bubble patterns as an analogy to the regular leaf arrangement in the context of phyllotaxis study (see, e.g., Ref. [26]). A phenomenological model was later developed by the same research group[27]. These two papers acknowledge the importance of bubble-bubble interactions and flow advection. Nevertheless, the way in which these factors contribute to pattern formation require further elucidation. In this paper, we revisit these regular surface patterns “drawn” by rising bubbles from a numerical perspective, aiming to gain a deeper understanding of the essential underlying physics of this self-organized phenomenon.
The rest of the paper is arranged as follows: Section II describes the numerical setup of the simulations. This is followed by section III, where the numerical results are presented and discussed. We further propose a heuristic model in section IV. A summary is given in section V.
@@ II Numerical setup
respectively, where $t$ is time, $\bm{u}$ is the velocity vector, $p$ is pressure, $\bm{g}$ is the gravitational acceleration whose magnitude is denoted by $g$, and $\bm{f}_{\sigma}$ represents the surface tension effect. $\rho$ and $\mu$ are the effective density and dynamic viscosity, respectively. They are calculated as weighted averages based on the local gas volume fraction $\alpha$:
@@
Eqs. (1) and (2) are solved using the second-order Bell-Colella-Glaz method[29]. The volume fraction field advection is handled using the volume-of-fluid (VOF) method, with a piecewise linear interface calculation (PLIC) method[30] employed for interface reconstruction. The surface tension term is defined as $\bm{f}_{\sigma}=\sigma\kappa\nabla\alpha$, where $\sigma$ denotes the surface tension coefficient. The interfacial curvature, $\kappa$, is evaluated using the method of particles[31].
Additionally, an improved multi-marker technique[28] is employed in the VOF method to prevent bubble coalescence, ensuring that bubbles can float on the free surface without bursting in the simulation. It is noteworthy that the present study focuses on cases in which bubble coalescence and bursting are absent. Under these conditions, the multi-marker method ensures fidelity in capturing bubble-bubble and bubble-free surface interactions without introducing non-physical artifacts[32, 28]. The simulation results and conclusions would not be valid if coalescence or bursting were to occur.
The accuracy of the solver is validated by simulating a single bubble rising in quiescent liquid. Four different liquids are tested with their physical parameters the same as those of cases S3, S5, S6, and S8 in Ref. [6]. The Morton numbers, defined as $\mathit{Mo}=g\mu_{l}^{4}/(\rho_{l}\sigma^{3})$, are 0.11, $9\times 10^{-4}$, $1\times 10^{-4}$, and $9\times 10^{-7}$ for these four cases, respectively. For each liquid, the volume-equivalent spherical diameter of the bubble, $d$, is varied across a relatively wide range. The computational domain is a liquid-filled box with a Cartesian mesh composed of cubic cells. The cell size is set to approximately $d/10$, maintaining consistent spatial resolution for different bubble sizes.
@@
The computational domain for simulations of the surface patterns is a square box partially filled with liquid, see Fig. 3(a). A Cartesian coordinate system is established with the origin $O$ located at the geometric center of the box’s bottom wall. The $z$-axis is oriented upward, while the $x$- and $y$-axes align with the side walls of the box. The dimensions of the box are $0.1,0.1]\,\text{m}\times[-0.1,0.1]\,\text{m}\times[0,0.075]\,\text{m}$, with the undisturbed free surface located at z=0.055z=0.055\,m. Gas bubbles with a volume-equivalent diameter of $d=4$ mm are released periodically from a source at (0, 0, 0.005)(0,\,0,\,0.005) m. We generate the bubbles by setting the gas volume fraction to 1 in a spherical region centered at the source. A uniform computational mesh with $N_{x}\times N_{y}\times N_{z}=512\times 512\times 192$ cubic cells is used, ensuring that there are about 10 mesh cells within the diameter of a spherical bubble. At this resolution, each case running in parallel on 128 CPU cores requires approximately four days. A finer mesh is also tested, yielding consistent results. More details can be found in Appendix A.
@@
Free-slip conditions are imposed at the top and bottom boundaries of the computational domain, while periodic conditions are applied on all side boundaries. In our simulations, the side boundaries are sufficiently far from the area of interest, namely, the central region of the free surface. Therefore, their influence on the bubble dynamics is negligible. It should be noted, however, that when side walls are close to the bubble chain, the vertical bubble motion can induce significant global recirculation in the confined space. This confined recirculation may subsequently alter the bubble velocity and, consequently, the final pattern formation. For the single-bubble cases shown in Fig. 2, a bubble acquires terminal velocity soon after release. However, because of the wake effects from preceding bubbles, for a bubble rising in a bubble chain, a much larger height is required for its velocity to reach a constant value. In the simulations of the current study, the bubbles are still gradually accelerating before approaching the free surface. Consequently, the liquid depth $h$ (or its dimensionless form $h/d$) is also a relevant parameter. Its effect will be briefly discussed in section V.
Figure 3: (a) Illustration of the computational domain (bubble size not to scale) and the coordinate system. (b) Simulated bubbles rising in a chain, with the free surface (black) and streamlines (gray) shown in a cross section through the $z$-axis. Only a small region close to the bubble chain is displayed.
In all simulations, the gravitational acceleration is set to $g=10\,\mathrm{m/s^{2}}$ in the z-z direction. The density and dynamic viscosity of the liquid ($l$) and gas ($g$) phases are: $\rho_{l}=1000\,\mathrm{kg/m^{3}}$, $\rho_{g}=30\,\mathrm{kg/m^{3}}$, $\mu_{l}=2\times 10^{-2}\,\mathrm{Pa\cdot s}$, and $\mu_{g}=2\times 10^{-5}\,\mathrm{Pa\cdot s}$. The surface tension coefficient $\sigma=0.02\,\mathrm{N/m}$. These physical parameters are similar to those used by Yoshikawa et al. [25], allowing a reasonable comparison between the numerical results and the experimental data in Ref. [25]. Given the low density ratio ($\rho_{g}/\rho_{l}=0.03$) and viscosity ratio ($\mu_{g}/\mu_{l}=0.001$) between gas and liquid, the effects of gas phase properties are expected to be minimal. The Ohnesorge and Bond numbers are calculated as $\mathit{Oh}=\mu_{l}/\sqrt{\rho_{l}\sigma d}\approx 0.071$ and $\mathit{Bo}=\rho_{l}gd^{2}/\sigma=8$, respectively. The corresponding Morton number is $\mathit{Mo}=\mathit{Oh}^{4}\mathit{Bo}=2\times 10^{-4}$, lying between the values of cases S5 and S6 in Fig. 2. The bubbles exhibit moderate deformation during ascent. No path instabilities are observed with the parameters investigated.
@@ III.1 Overview
With the VOF method, the motion and deformation of each bubble are fully resolved. Figure 3(b) shows a cross-sectional view of the numerical result for a typical case we have simulated. The bubbles rise along a straight path in a chain and are arrested by the free surface. As can be seen from the streamlines, the bubble chain induces a vertical liquid jet along the central axis. This jet later on converts into a divergent radial flow in a thin layer beneath the free surface. Owing to the overshooting effect of the bubbles and jet, a small circular bump forms on the free surface, centered at the bubble emergence site. These observations agree well with those reported in Ref. [27].
Systematic simulations are performed to examine the effect of the bubble release period $T$ on pattern formation, with other parameters held constant. The top panels of Fig. 4 depict four representative surface patterns observed in the simulations, with $T$ decreasing from left to right. Panel (a) shows a disordered mode, characterized by a random distribution of bubbles on the free surface. Panel (b) displays a two-armed mode, where the bubbles align along two branches in a straight line. At smaller bubble release periods, multi-armed modes can be observed, as shown in panels (c) and (d). The corresponding movies for all these four cases can be found in the Supplemental Material[33].
Figure 4: (a) to (d): Four typical surface patterns observed in the numerical simulations (top view). The free surface is colored in gray. The dashed lines serve as visual guides to distinguish between different branches. The numbers indicate the sequence in which successive bubbles join these branches. From left to right, the release periods are 100, 50, 42, and 36 ms, respectively. (e) to (h): Divergence angle as a function of bubble release sequence number for the corresponding cases in the top row. The red dashed lines show the averaged values.
It is known that the pattern formation is closely related to the directions in which the bubbles are emitted in sequence. In the bottom panels of Fig. 4, the difference $\varphi$ between the emission directions of consecutive bubbles, referred to as the “divergence angle” in Ref. [25], is shown as a function of bubble release sequence number. As will be subsequently elucidated, $\varphi$ is fully determined by the local dynamics in the vicinity of the bubble emergence site. In detail, four major processes are identified to be responsible for the formation of different surface patterns. They are (i) pair collisions, (ii) non-contact repulsion, (iii) bump confinement, and (iv) flow advection. The first two processes both describe the interaction between bubbles. Pair collision occurs when two bubbles are close enough with their mass center distance $\ell\lesssim d$, and is associated with significant bubble squeezing, deformation, and quick velocity changes. Whereas non-contact repulsion is effective when $\ell\gtrsim d$ and is less violent.
@@ III.2 Disordered and fixed two-armed modes
The disordered mode is seemingly the most common situation. It occurs when the bubble release period $T$ is relatively large ($T\gtrsim 100\,\mathrm{ms}$ in the case illustrated in Fig. 4). This mode is characterized by random pair collisions between bubbles. For instance, as bubble 1-1 rises to the free surface, it collides with bubble 2-2 which is already at the emergence site. The two bubbles are knocked away by each other in opposite directions, creating a $180^{\circ}$ divergence angle. By the time bubble 0 reaches the surface, the influence of the previous collision has vanished. Bubble 0 stays at the center and collides with the subsequent bubble 1, resulting in opposite emission directions that do not necessarily align with those of bubbles 2-2 and 1-1. As a consequence, the divergence angle jumps between $180^{\circ}$ and other random values (Fig. 4e). Additionally, a large $T$ results in a relatively weak liquid jet, such that the bubbles tend to aggregate near the central region of the free surface. Both of these features hinder the formation of regular patterns with distinguishable numbers of arms.
The reason for the transition from the disordered mode to a fixed two-armed mode (Fig. 4b) is straightforward: When $T$ is decreased below a certain threshold value, bubble 1-1, after collision with bubble 2-2, does not have enough time to move far away from the center before bubble 0 rises to the free surface. Consequently, bubble 0 collides with bubble 1-1 and bounces towards the opposite direction to the latter’s position vector. Similarly, bubble 1 is emitted in the opposite direction to that of bubble 0, following the path of bubble 1-1. In this way, the bubbles alternatively join two branches with a constant divergence angle of $\varphi=180^{\circ}$, as depicted in Fig. 4(f). By conducting simulations with gradually decreased bubble release period, for the case shown in Fig. 4, this transition is found to occur at $T\approx 80\,\mathrm{ms}$.
@@ III.3 Multi-armed modes
The multi-armed modes occur when the bubble release period is further decreased. They exhibit richer phenomena compared with the disordered and fixed two-armed situations. Figures 4(c) and 4(d) show a fixed five-armed pattern ($T=42$ ms) and a spiral three-armed pattern ($T=36$ ms), respectively. A 3D view of a single bubble trajectory in the five-armed case is displayed in Fig. 5(a). The projection of this trajectory onto the horizontal plane is shown in Fig. 5(b) alongside the trajectories of two consecutive bubbles. An amplified view is provided in the inset, highlighting the curved trajectories of the bubbles near the emergence site.
Figure 5: (a) Free surface (gray) and trajectory (cyan) of a single bubble in a five-armed case. (b) Trajectories of three consecutive bubbles projected onto the horizontal plane. The green dashed lines indicate the emission directions of bubbles 2-2 and 1-1. The inset is an amplified view of the trajectories in the vicinity of the bubble emergence site.
@@
The twisted trajectories in Fig. 5 indicate complex bubble dynamics in the vicinity of the bubble emergence site. Yoshikawa et al. [25] note that, despite the arms may rotate, “the motion of each bubble is purely radial.” This is true except for in the short-period initial phase of the bubble motion, as can be seen in Fig. 5(b). To elucidate the entire process governing the divergence angle, we divide the trajectory of a single bubble into six distinct stages, labeled I to VI, as indicated above the trajectory plot in Fig. 6. This plot displays the bubble’s path projected onto the horizontal plane, with the black dots marking positions at equal time intervals (the sparseness of the dots is thus an indicator of the bubble’s horizontal-velocity magnitude). The accompanying insets provide close-up views of the instantaneous configuration of the bubbles near the free surface. The dynamics of the bubbles and their connection with the processes listed in section III.1 are described in detail below.
@@
Firstly, a significant difference between multi-armed modes and the simpler fixed two-armed mode is that the former cannot be explained solely by binary bubble collisions. This adds complexity to the analysis of the underlying physics, necessitating the inclusion of process (ii), i.e., non-contact repulsion. This process is illustrated in Fig. 7, where three snapshots of the three-armed situation are shown in chronological order. During this stage, bubble 0 is still rising below the free surface without colliding with preceding bubbles. Floating on the free surface with a short distance from the center where the bubbles emerge, bubble 2-2 acts as an obstacle that partially blocks the rightward motion of the liquid flow. As a consequence of this asymmetry, bubble 0’s rising path is deflected to the left, as if it experiences repulsion from bubble 2-2. Since there is no direct contact between bubbles 0 and 2-2 throughout the interaction, the horizontal displacement of bubble 0 from the central axis is mild. We denote this displacement by $\bm{\lambda}_{1}$. This process corresponds to stage I in Fig. 6(a). Similar interactions also occur between bubbles 0 and 1-1. However, since bubble 1-1 is only slightly deviated from the center, its influence on bubble 0’s horizontal motion is much weaker than the influence of bubble 2-2. The repulsion from other floating bubbles is negligible due to their larger distance from bubble 0. Therefore, $\bm{\lambda}_{1}$ is roughly in the opposite direction of bubble 2-2’s position vector in the horizontal plane, as shown in the inset of Fig. 5(b).
@@
The motion of bubble 0 in stage II is governed by process (i), i.e., pair collision. Figure 8 depicts the collision process between bubbles 0 and 1-1 in the cross-section where the centers of the two bubbles are located. The snapshot in Fig. 8(a) is taken at about the same time as that in Fig. 7(c). When approaching the free surface, bubble 0 is impeded by bubble 1-1, squeezed and pushed away by it. During this process, bubble 0 acquires a horizontal displacement roughly in the opposite direction of bubble 1-1’s position vector, which can also be observed in the inset of Fig. 5(b). We denote this displacement by $\bm{\lambda}_{2}$.
@@
Although processes (i) and (ii) both have effects to accelerate bubble 0, the bubble’s radial velocity falls back to nearly zero by the end of stage II. This can be seen from Fig. 9, where the radial velocities of bubbles 2-2, 1-1, 0, and 1 are shown as functions of time. This deceleration of bubble 0 is due to process (iii) mentioned in section III.1, i.e., the confinement of the bump. As revealed in Fig. 8(c), after collision with bubble 1-1, bubble 0 is located in the very top region of the bump. Thus, its horizontal motion is restricted by the capillary effect of the deformed and more vertical side surface of the bump.
Figure 9: Radial velocity versus time for bubbles 2-2, 1-1, 0, and 1 (from left to right). Stages are labeled with respect to bubble 0.
It is noteworthy that, up to this point, bubble 0’s displacement from the central axis is no greater than half its radius. However, the final emission direction (the red dash-dotted line in Fig. 6a) of the bubble has been essentially determined. All the remaining stages play a role to “launch” this nearly stationary bubble and transport it in the same direction as its current position vector.
The collision with bubble 0 also promotes bubble 1-1 to leave the central region. As a consequence, bubble 0 has more room to flatten, and the bump surface becomes less steep. This relieves the constraint imposed by the bump, allowing bubble 0 to regain a small velocity away from the center advected by the radial flow mentioned earlier. This process corresponds to stage III in Figs. 6 and 9.
@@
Being pushed away from the center by bubble 1, the subsequent motion of bubble 0 in stage V is driven by the divergent radial flow (process (iv) in section III.1). This effect of flow advection is demonstrated by the streamlines shown in Fig. 10. The bubble continues moving within the bump until it encounters the edge of it, where the bubble experiences a temporary deceleration due to the transition between the bumped and horizontal free surface. This can be seen in Fig. 11, where the vertical dashed line indicates that the local minimum radial velocity occurs exactly when the bubble reaches the flat liquid surface from a higher elevation.